# Aubry-André Transition for an Active Undulator

### v1

@@
Aubry-André Transition for an Active Undulator
Christopher J. Pierce, Tianyu Wang, Dmitri Kalinin,
Andrew Zangwill and Daniel I. Goldman
School of Physics, Georgia Institute of Technology
Authors contributed equally
To whom correspondence should be addressed
@@
The transport of deformable self-propelling objects like bacteria, worms, snakes, and robots through heterogeneous environments is poorly understood. In this paper, we use experiment, simulation, and theory to study a snake-like robot as it undulates without sensory feedback through a narrow channel containing a linear array of boulder-like hemispherical obstacles. The profile of the boulder landscape approximates a one-dimensional bichromatic potential introduced by Aubry and André (AA) to study the propagation and localization of waves. When the boulder landscape is strictly periodic, the robot can pass completely through the channel. But if the landscape is sufficiently aperiodic, the robot can become trapped and fail to exit the channel. The metrics we use to quantify this change agree well with earlier experimental and theoretical work on a localization transition that occurs when quantum waves interact with the AA potential. A theoretical treatment of the robot’s motion using resistive force theory modified to include spatially varying drag forces reproduces the behavior we observe.
Introduction
@@
It is well known that equations of motion derived for point particles subject to simple external forces make predictions that agree well with experimental observations, e.g., electrons drifting through a current-carrying wire, biomolecules diffusing through a gel, and asteroids hurtling through space. Compared to these systems, much less attention has been paid to systems that actively self-deform and self-propel, e.g., bacteria that swim through fluids by rotating their flagella  (?, ?), snakes that move through sand by bending their bodies  (?, ?, ?) and robots that use legs to walk on granular media (?). For locomotors of this kind, forward motion results from environmental reaction forces such as friction and fluid drag. These forces typically vary over both the length of the object and in time because they arise in response to complex, internally powered and controlled body deformations. Unlike a point particle, the principles that govern the transport of a self-propelled and self-deforming object must reflect both the geometry of its self-deformations—known as gaits—and the detailed nature of the reaction forces exerted on it by its environment.
@@
Despite these complications, researchers from various disciplines have developed useful descriptions of the self-propulsion of objects through environments that are homogeneous and isotropic  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?). The same cannot be said for environments that are inhomogeneous in space, time, or both. In the living world, this is the situation faced by spermatozoa swimming in complex heterogeneous fluids, snakes slithering through a garden, and horses competing in a steeplechase. In the human-made world, familiar examples include autonomous search-and-rescue robots  (?), extraterrestrial rovers  (?, ?), and agricultural robots design to survey and protect crops  (?, ?).
@@
Previous work revealed that snakes, nematode worms, and an elongated limbless robot successfully traverse certain periodically heterogeneous environments by exploiting a strategy they typically use to traverse homogeneous environments. That is, they impose a quasi-sinusoidal bending wave over the length of their bodies  (?, ?, ?, ?, ?, ?). In one study, snakes and an undulating robot deforming in this way passed through a one-dimensional array of vertical posts in a manner reminiscent of the way light diffracts from an array of equally spaced slits cut into an opaque screen (?, ?). In another study, nematode worms and a limbless snake-like robot passed through a two-dimensional array of vertical posts at constant speed by mimicking the Bragg condition of crystallography and matching (approximately) their undulation wavelength to the periodicity of the array (?).
@@
Our snake-like robot was constructed from a set of $N=9$ servomotors (Dynamixel XC330-M288-T) connected by hinged joints (Fig. 1A). To mimic the undulatory gaits observed for certain snakes, nematode worms, and sandfish lizards  (?), we imposed a “serpenoid”   (?) deformation pattern on the robot’s body (Fig. 1B, C) by specifying that the joint angles vary in time as (Fig 1B)
@@
This produces an approximately sinusoidal bending wave of lateral displacement that travels from the robot’s head to its tail over one period. [^1] Because the number of joints is finite, it is necessary to multiply the factor of $2\pi m/N$ in Eq. (1) by a factor of $(N+1)/N$. We used $A=50^{\circ}$ and $\omega=0.0625$ ${\rm s}^{-1}$ for all the work reported in this paper. The robot’s joints are actuated by servomotors operating under position control. Each motor includes a built-in low-level PD controller, with default proportional and derivative gains set to $K_{P}=8.5$ and $K_{D}=31.2$, respectively. To complement our experiments, we developed a multibody simulation, [^2] Using CoppeliaSim  (?), which replicates the mechanical properties and control strategies of the physical robot. These simulations reproduce the behavior of the real robot well and therefore allowed us to systematically explore a broader range of parameter variation on a scale impractical for physical trials alone. As in our previous experiments  (?), we operated the robot in a regime where coasting  (?) is negligible. That it, self-propulsion ceases almost immediately when the deformations cease because frictional dissipation dominates inertia  (?).
We limited our robot to quasi-one-dimensional motion by confining it to a straight channel formed by two parallel rigid walls. Inside the channel, we arranged obstacles in a pattern inspired by a potential introduced by Aubry and André (AA) in their study of electrons in one-dimensional crystals  (?). In their work, electrons interact with a base potential $\propto\sin(kx)$ and a perturbation potential $\propto\sin(\beta kx)$. The eigensolutions associated with the AA potential change character from propagating to localized as a function of the value of the commensurability parameter $\beta$ and the relative amplitude of the two components of the potential  (?, ?).
@@
In detail, we first placed in the channel a row of equally-spaced 3D-printed spheres (diameter $d=6.3$ cm) truncated so their tops lay a distance $s=0.2d$ above the channel floor (top panel of Fig. 1C). [^3] This was the smallest value of $s$ where the robot moved one body length per one cycle of the imposed deformation wave. The separation $a$ between neighboring spheres (hereafter “boulders”) was chosen as one-half the wavelength of the deformation wave imposed on the robot. This choice facilitates the transmission of the robot through the array (?). To this periodic array of boulders, we added a perturbation in the form of a second row of equally spaced boulders with nearest-neighbor separation $a/\beta$ and truncated so their tops lay a height $h$ above the channel floor. For convenience, we will use the term “AA terrain” when referring to this obstacle course inside the channel. Our experiments and simulations consisted of initiating the robot at one end of the channel and measuring its body shape and center-of-mass position as a function of time for different values of $\beta$ and different heights of the perturbing boulder array between $h=0$ (fully buried) to $h=d$ (fully exposed).
@@
For many particle transport problems, it is possible to identify dynamical transitions from the fact that the mean square displacement (MSD) varies at long times as $\langle X^{2}(t)\rangle\sim at^{b}$. For example, Mokhtari and Zippelius (?) simulated the motion of active filaments through a porous medium and discovered a transition from ballistic ($b=2$) to diffusive ($b=1$) to localized ($b=0$) behavior as a function of the stiffness of the filament. In a similar way, the MSD exponent we calculate from our robot data (experiment and simulation) demonstrates that localization replaces ballistic motion when $h/d$ exceed a critical value (Fig. 2C).
Statistical distributions of the total travel distance $D$ provide another quantitative measure of transport change. We plot these distributions as solid curves in Fig. 3A for the same values of $h/d$ used in Fig. 2A. This illustrates the transport behavior before (green), during (blue), and after (red) the transition. Fitting these distributions to the function $Ae^{-(D-D_{0})^{\alpha}}$ (dashed curves in Fig. 3A) reveals a transition in the distribution shape from Gaussian ($\alpha=2$) to exponential ($\alpha=1$) (Fig. 3B). Remarkably, curves very similar to Fig. 2B and 3B appear in the work of Roati and co-workers (?, ?) who followed the motion of a Bose-Einstein condensate of ${\rm{}^{39}K}$ atoms through a one-dimensional optical potential of the AA type and observed the predicted transition from a de-localized quantum state to a localized quantum state.
@@
No single theory models the dynamics of an extended self-deforming object in all circumstances. However, when such an object moves through a homogeneous overdamped environment, so-called resistive force theory (RFT)  (?, ?) provides a good description. In this section, we describe a generalization of this theory capable of describing the undulatory transport of a snake-like robot through an inhomogeneous environment.
Following Gray and Hancock  (?), we replace the robot by a flexible one-dimensional filament of length $L$ on the $x$-axis. We model the effect of the robot’s internally generated bending wave by specifying the transverse displacements of the filament:
@@
The center-of-mass of the filament (hereafter “undulator”) moves along the $x$ axis because the bending described by Eq. (2) induces time-varying contact and resistive forces. The theory is designed for overdamped, low-coasting situations where inertia can be neglected. Accordingly, Gray and Hancock derived an expression for the undulator’s center-of-mass velocity, $V_{\rm cm}$, by setting to zero the sum of all the forces that act upon it. Their expression has been used with success to describe the transport of real undulators over flat frictional surfaces (?) and through viscous fluids (?) and frictional fluids like sand (?).
The key parameter in RFT is a single number called the drag anisotropy. This quantity, $\xi=f_{\perp}/f_{\parallel}$, is the ratio of the components of the total force that act perpendicular and parallel to the surface of the undulator. When $\xi>1$, the center-of-mass moves in the direction opposite to the direction of the propagating bending wave. This is the case for snakes on sand (?, ?), limbless robots with passive wheels on hard ground (?), and nematode worms (?) or sperm cells in low Reynolds number fluids (?), By contrast, the center-of-mass moves in the same direction as the deformation wave for systems with $\xi<1$ like centipedes on water surfaces (?) and some marine worms in bulk fluids  (?).
A simple way to generalize RFT for use with inhomogeneous terrains is to permit the drag anisotropy $\xi$ to vary as a function of position. [^4] The numerical studies of limbless locomotion by Zhang and co-workers  (?) introduce a spatially-varying planar friction coefficient to study the effects of terrain inhomogeneity. An important feature of this approach is that the same body deformation wave can induce either forward or backward motion depending on the geometry of the contact between the undulator and the terrain. To illustrate how this can happen, the side and top views in Fig. 4A show the normal ($N$), frictional ($F$), and gravitational ($mg$) forces present when a thin rod strikes a hemispherical boulder tangentially. The magnitude and direction of the net force $\vec{f}$ on the rod depend sensitively on the exact point of contact with the boulder. For example, when the rod is moving to the right, it is straightforward to confirm that $\xi>1$ on the left side of the boulder (moving “uphill”) while $\xi<1$ on the right side of the boulder (moving “downhill”).
We have not attempted to calculate the spatial variation of the drag anisotropy for any actual terrain. Instead, we construct a minimal model that uses simple analytic anisotropy functions $\xi(x)$ that possess the same spatial periodicities as the AA terrain. With $\xi(x)$ in hand, the generalization of the expression for $V_{\rm cm}$ derived in Ref.  (?) for small amplitude undulations can be shown to be (see the Supplemental Material)
@@
Fig. 5A plots our RFT results for the average distance traveled by the undulator as a function of $\Gamma/\Delta$ for three values of $\beta$. The good qualitative agreement between the $\beta=(\sqrt{5}-1)/2$ curve (red) and the corresponding experimental data plotted in Fig. 5B (repeated from Fig. 2B) gives us further confidence that RFT captures the essential dynamics of the real robot moving through an AA terrain. The RFT results shown in Fig. 5A for rational values of $\beta$ are surprising (initially) because one reflects ballistic motion ($\beta=1/2$ in green) and the other exhibits localization ($\beta=2/3$ in blue). However, comparison with Fig 5B shows that the robot behaves in exactly the same way. Moreover, the qualitative difference in transport behavior associated with these two rational values of $\beta$ is precisely what is seen in detailed numerical studies of the AA model where the eigenstates of the Schrödinger equation change from delocalized to localized even when $\beta=p/q$ as long as $p\neq 1$ and $q$ are integers (?).
@@
Fig. 6D plots $P_{\rm broad}/P_{\rm 0}$ for the robot calculated similarly using simulation data along with the probability that a trial ends in termination (redrawn from Fig. 2B). Once again, the torque noise rises and falls as a function of the terrain aperiodicity, albeit more steeply in Fig. 6D than in Fig. 6C. The reason for this, we posit, is that the three-dimensional robot can respond to aperiodicity in many ways while the RFT undulator, which is constrained by construction to 1-D, can do so only by reversals and oscillatory trapping. For example, when $h/d$ is small (below the peak in Fig. 6D), the robot can negotiate the terrain using small out-of-plane motions which produce less broad-band torque noise than a reversal event. As the aperiodicity becomes higher, eventually the robot fails to respond with small out-of-plane motion, and begins to pick up noise in the torque signal. When $h/d$ is large (above the peak), the robot can flip over, jam against a sidewall, or distort into a shape very far from serpenoid. All of the latter events trigger termination, which leads us to conclude that termination simply preempts localization by trapping and likely occurs in response to the same torque patterns. That is why, despite the somewhat ad hoc rules we developed to initiate termination, the systematics of the transition to localization exhibited by our experiments and simulations qualitatively agree with both our RFT calculations and with the behavior expected from the quantum AA model.
@@
The similarity of our results to the cold atom experiments deserves further comment. We are aware that a microscopic quantum particle like an atom and a macroscopic self-deforming object like our robot have nothing in common except that quantum mechanics requires the atom to possess a de Broglie wave, and the act of moving requires our robot to possess a deformation wave. That being said, it is interesting to note  (?) that a transition from a delocalized state to a localized state is possible for both classical and quantum particles when they interact in one dimension with the AA potential. Moreover, not unlike our RFT undulator, localization of these particles takes the form of trapping by the wells of the AA potential. The same is not true for the one-dimensional Anderson model, where the corresponding potential is random and delocalization can occur for a classical particle but never for a quantum particle.
In closing, we suggest it is worth considering our present results in light of earlier experiments where we observed diffraction-like behavior when snakes and limbless robots passed through a one-dimensional array of posts. We also note work by Zhang and co-workers (?) who found lensing, refraction, and wave-guiding behavior in numerical studies of a filamentary undulator sliding along a heterogeneous surface tailored to mimic optical elements. Taken together, it is tempting to speculate that there is some previously unsuspected connection between the waves associated with self-deforming mechanical systems and the more familiar waves studied in classical and quantum mechanics.
Acknowledgments
This work was funded by NSF 2310751 and DOD W911NF2110033.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S3
Tables S1 to S4
References (4-10)
Supplementary Materials for
Transport and Localization of a Self-Deforming Body through a Disordered Medium
Christopher J. Pierce1∗, Tianyu Wang1∗, Dmitri Kalinin1,
Andrew Zangwill1 and Daniel I. Goldman1,†
1School of Physics, Georgia Institute of Technology
Authors contributed equally
To whom correspondence should be addressed
The PDF file includes: Materials and Methods
Supplementary Text
Figs. S1 to S3
Tables S1 to S4
References (4-10)
Materials and Methods
@@ Acknowledgments
To complement our experimental investigation of undulatory transport in structured environments, we developed a multibody dynamics simulation using the CoppeliaSim platform  (?). The simulation replicates the kinematics of the undulating robot by prescribing a traveling wave of joint angles consistent with a serpenoid gait. While the number of body segments in the simulation was chosen to support computational efficiency and broader parametric exploration, the wavelength of body undulation was kept consistent with that used in physical experiments to preserve comparable locomotor-environment interaction dynamics.
CoppeliaSim was used in conjunction with its Bullet physics  (?) backend. Contact interactions were resolved using Bullet’s default sequential impulse constraint solver, an iterative method that applies impulses to enforce non-penetration and friction constraints. This constraint-based formulation enables stable simulation of rigid body interactions in friction-dominated regimes, even in highly cluttered environments.
@@
The robot in simulation was constructed as a planar multibody system using the exact geometry of the physical robot, imported directly from computer-aided design models built in SolidWorks. This ensured that segment dimensions, mass distribution, and contact surface geometry were faithfully preserved, enabling direct comparison between simulated and experimental results. Adjacent segments were connected via revolute joints allowing relative yaw motion. Joint actuation followed a serpenoid waveform of the form:
$\alpha_{m}(t)$ $=A\sin(2\pi m/N-\omega t),\hskip 28.45274ptm=1,\dots N$ (S1)
@@
Simulations were terminated early if the robot flipped out of the plane, defined as any point on the robot body exceeding a fixed vertical threshold (15 cm, ${}\sim 0.9$ width of the body wave) in the $z$-direction. No restitution or damping was applied, allowing motion to cease naturally under frictional contact.
@@
$\frac{dF}{dx}=(C_{N}-C_{L})\frac{dy}{dx}\frac{dy}{dt}-C_{L}\frac{dX}{dt}.$ (S3)
The variable $x\in[0,L]$ describes the position along the body in a coordinate system that moves with a point at the tail, while the variable $X(x,t)$ describes the position of a point $x$ along the body in a fixed reference frame as moves. The variables $C_{N}$ and $C_{L}$ refer to the normal and longitudinal resistance coefficients (in this case, viscous drag coefficients), respectively, which were assumed by Gray and Hancock to be constants intrinsic to the material constituting the body and the environment. In this manuscript we will eventually allow these constants to vary in space (along $x^{\prime}$), to describe locomotion in heterogeneous environments. We note that this relation is valid in the limit of small displacements $y$.
@@
Substituting (S4) into (S3) yields
$\frac{dF}{dx}=4\pi^{2}y_{0}^{2}(C_{N}-C_{L})\cos^{2}{2\pi(x+t)}-C_{L}\frac{dX}{dt}$ (S5)
@@
We assume that each segment of the body moves along the X-axis at the same speed (the body does not deform along the swimming direction, but only lateral to the swimming direction), such that integrand on the lhs of (S6) is a constant. We now introduce the spatial dependence of the resistance coefficients by writing them as follows
$\frac{C_{N}-C_{L}}{C_{L}}=\frac{C_{N}(X)}{C_{L}(X)}-1=\xi(X)-1$ (S7)

### v2

# pnasresearcharticle

@@
Abstract
@@
The transport of deformable self-propelling objects like bacteria, worms, snakes, and robots through heterogeneous environments is poorly understood. In this paper, we use experiment, simulation, and theory to study a snake-like robot as it undulates without sensory feedback through a narrow channel containing a linear array of boulder-like hemispherical obstacles. The profile of the boulder landscape approximates a one-dimensional potential introduced by Aubry and André (AA) to study wave function localization in aperiodic lattices. The AA model provides a deterministically disordered alternative to the better-known phenomenon of Anderson localization, which occurs in truly random disordered lattices. When the boulder landscape is strictly periodic, the robot can pass completely through the channel. But if the landscape is sufficiently aperiodic, the robot becomes trapped and fails to exit the channel. The metrics we use to quantify this transition – including exponential distributions of robot position when localized – agree well with earlier experimental and theoretical work on a localization transition that occurs when quantum waves interact with the AA potential. A theoretical treatment of the robot’s motion using resistive force theory modified to include spatially varying drag forces reproduces the behavior we observe. Further, our results indicate that the transition is generated by large fluctuations in the driving torques required for self-propulsion. These results point to a potentially fundamental connection between classical and quantum wave mechanics and the locomotion of undulators. Our study illustrates how analogies with models from condensed matter physics and wave optics can lead to the discovery of principles of self-propulsion in non-periodic landscapes.
keywords:
Undulatory Locomotion 1 |||| Localization 2 |||| Disordered Media 3 ||||
[…]
This manuscript was compiled on July 23, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
8]3
[…]
@@
Equations of motion derived for point particles subject to simple external forces make predictions that agree well with experimental observations, e.g., electrons drifting through a current-carrying wire, biomolecules diffusing through a gel, and asteroids hurtling through space. Compared to these systems, much less attention has been paid to systems that actively self-deform and self-propel, e.g., bacteria that swim through fluids by rotating their flagella (1, 2), snakes that move through sand by bending their bodies (3, 4, 5) and robots that use legs to walk on granular media(6). For locomotors of this kind, forward motion results from environmental reaction forces such as friction and fluid drag. These forces typically vary over both the length of the object and in time because they arise in response to complex, internally powered and controlled body deformations. Unlike a point particle, the principles that govern the transport of a self-propelled and self-deforming object must reflect both the geometry of its self-deformations—known as gaits—and the detailed nature of the reaction forces exerted on it by its environment.
@@
Despite these complications, researchers from various disciplines have developed useful descriptions of the self-propulsion of objects through environments that are homogeneous and isotropic (7, 8, 1, 9, 3, 10, 11, 4, 2, 12). The same cannot be said for environments that are inhomogeneous in space, time, or both. In the living world, this is the situation faced by spermatozoa swimming in complex heterogeneous fluids, snakes slithering through a garden, and horses competing in a steeplechase. In the human-made world, familiar examples include autonomous search-and-rescue robots (13), extraterrestrial rovers (14, 15), and agricultural robots design to survey and protect crops (16, 17).
@@
Previous work revealed that snakes, nematode worms, and an elongated limbless robot successfully traverse certain periodically heterogeneous environments by exploiting a strategy they typically use to traverse homogeneous environments. That is, they impose a quasi-sinusoidal bending wave over the length of their bodies (18, 19, 20, 21, 22, 23, 24). In one study, snakes and an undulating robot deforming in this way passed through a one-dimensional array of vertical posts in a manner reminiscent of the way light diffracts from an array of equally spaced slits cut into an opaque screen(21, 20). In another study, nematode worms and a limbless snake-like robot passed through a two-dimensional array of vertical posts at constant speed by mimicking the Bragg condition of crystallography and matching (approximately) their undulation wavelength to the periodicity of the array(22).
@@ 1 Results and Discussion
Our snake-like robot was constructed from a set of $N=9$ servomotors (Dynamixel XC330-M288-T) connected by hinged joints (Fig. 1A). To mimic the undulatory gaits observed for certain snakes, nematode worms, and sandfish lizards (12), we imposed a “serpenoid”  (25) deformation pattern on the robot’s body (Fig. 1B, C) by specifying that the joint angles vary in time as (Fig 1B)
@@
This produces an approximately sinusoidal bending wave of lateral displacement that travels from the robot’s head to its tail over one period. [^1] Because the number of joints is finite, it is necessary to multiply the factor of $2\pi m/N$ in Eq. (1) by a factor of $(N+1)/N$. We used $A=50^{\circ}$ and $\omega=0.0625$ ${\rm s}^{-1}$ for all the work reported in this paper. The robot’s joints are actuated by servomotors operating under position control. Each motor includes a built-in low-level PD controller, with default proportional and derivative gains set to $K_{P}=8.5$ and $K_{D}=31.2$, respectively. To complement our experiments, we developed a multibody simulation, [^2] Using CoppeliaSim (26), which replicates the mechanical properties and control strategies of the physical robot. These simulations reproduce the behavior of the real robot well and therefore allowed us to systematically explore a broader range of parameter variation on a scale impractical for physical trials alone. As in our previous experiments (22), we operated the robot in a regime where coasting (12) is negligible. That it, self-propulsion ceases almost immediately when the deformations cease because frictional dissipation dominates inertia (27).
We limited our robot to quasi-one-dimensional motion by confining it to a straight channel formed by two parallel rigid walls. Inside the channel, we arranged obstacles in a pattern inspired by a potential introduced by Aubry and André (AA) to study the localization of electrons in one-dimensional aperiodic lattices (28); for a useful pedagogical discussion of the AA model, see (29). This potential can be thought of as a pseudo-randomly disordered version of the truly randomly disordered potential used by Anderson in his pioneering study of electron localization.(30) In the AA model, electrons interact with a base potential $\propto\sin(kx)$ and a perturbation potential $\propto\sin(\beta kx)$. The eigensolutions associated with the AA potential change character from propagating to localized as a function of the value of the commensurability parameter $\beta$ and the relative amplitude of the two components of the potential (31, 32).
@@
In detail, we first placed in the channel a row of equally-spaced 3D-printed spheres (diameter $d=6.3$ cm) truncated so their tops lay a distance $s=0.2d$ above the channel floor (top panel of Fig. 1C). [^3] This was the smallest value of $s$ where the robot moved one body length per one cycle of the imposed deformation wave. The separation $a$ between neighboring spheres (hereafter “boulders”) was chosen as one-half the wavelength of the deformation wave imposed on the robot. This choice facilitates the transmission of the robot through the array(22). To this periodic array of boulders, we added a perturbation in the form of a second row of equally spaced boulders with nearest-neighbor separation $a/\beta$ and truncated so their tops lay a height $h$ above the channel floor. For convenience, we will use the term “AA terrain” when referring to this obstacle course inside the channel. Our experiments and simulations consisted of placing the robot at one end of the channel, initiating the gait sequence, and measuring its body shape and center-of-mass position as a function of time for different values of $\beta$ and different heights of the perturbing boulder array between $h=0$ (fully buried) to $h=d$ (fully exposed).
@@
For many particle transport problems, it is possible to identify dynamical transitions from the fact that the mean square displacement (MSD) varies at long times as $\langle X^{2}(t)\rangle\sim at^{b}$. For example, Mokhtari and Zippelius(33) simulated the motion of active filaments through a porous medium and discovered a transition from ballistic ($b=2$) to diffusive ($b=1$) to localized ($b=0$) behavior as a function of the stiffness of the filament. In a similar way, the MSD exponent we calculate from our robot data (experiment and simulation) demonstrates that localization replaces ballistic motion when $h/d$ exceed a critical value (Fig. 2C).
Statistical distributions of the total travel distance $D$ provide another quantitative measure of transport change. We plot these distributions as solid curves in Fig. 3A for the same values of $h/d$ used in Fig. 2A. This illustrates the transport behavior before (green), during (blue), and after (red) the transition. Fitting these distributions to the function $Ae^{-(D-D_{0})^{\alpha}}$ (dashed curves in Fig. 3A) reveals a transition in the distribution shape from Gaussian ($\alpha=2$) to exponential ($\alpha=1$) (Fig. 3B). Remarkably, curves very similar to Fig. 2B and 3B appear in the work of Roati and co-workers(34, 35) who followed the motion of a Bose-Einstein condensate of ${\rm{}^{39}K}$ atoms through a one-dimensional optical potential of the AA type and observed the predicted transition from a de-localized quantum state to a localized quantum state.
@@
No single theory models the dynamics of an extended self-deforming object in all circumstances. However, when such an object moves through a homogeneous overdamped environment, so-called resistive force theory (RFT) (8, 11) provides a good description. In this section, we describe a generalization of this theory capable of describing the undulatory transport of a snake-like robot through an inhomogeneous environment.
Following Gray and Hancock (8), we replace the robot by a flexible one-dimensional filament of length $L$ on the $x$-axis. We model the effect of the robot’s internally generated bending wave by specifying the transverse displacements of the filament:
@@
The center-of-mass of the filament (hereafter “undulator”) moves along the $x$ axis because the bending described by Eq. (2) induces time-varying contact and resistive forces. The theory is designed for overdamped, low-coasting situations where inertia can be neglected. Accordingly, Gray and Hancock derived an expression for the undulator’s center-of-mass velocity, $V_{\rm cm}$, by setting to zero the sum of all the forces that act upon it. Their expression has been used with success to describe the transport of real undulators over flat frictional surfaces(20) and through viscous fluids(8) and frictional fluids like sand(11).
The key parameter in RFT is a single number called the drag anisotropy. This quantity, $\xi=f_{\perp}/f_{\parallel}$, is the ratio of the components of the total force that act perpendicular and parallel to the surface of the undulator. When $\xi>1$, the center-of-mass moves in the direction opposite to the direction of the propagating bending wave. This is the case for snakes on sand(36, 37), limbless robots with passive wheels on hard ground(20), and nematode worms(38) or sperm cells in low Reynolds number fluids(8), By contrast, the center-of-mass moves in the same direction as the deformation wave for systems with $\xi<1$ like centipedes on water surfaces(39) and some marine worms in bulk fluids (7).
A simple way to generalize RFT for use with inhomogeneous terrains is to permit the drag anisotropy $\xi$ to vary as a function of position. [^4] The numerical studies of limbless locomotion by Zhang and co-workers (40) introduce a spatially-varying planar friction coefficient to study the effects of terrain inhomogeneity. An important feature of this approach is that the same body deformation wave can induce either forward or backward motion depending on the geometry of the contact between the undulator and the terrain. To illustrate how this can happen, the side and top views in Fig. 4A show the normal ($N$), frictional ($F$), and gravitational ($mg$) forces present when a thin rod strikes a hemispherical boulder tangentially. The magnitude and direction of the net force $\vec{f}$ on the rod depend sensitively on the exact point of contact with the boulder. For example, when the rod is moving to the right, it is straightforward to confirm that $\xi>1$ on the left side of the boulder (moving “uphill”) while $\xi<1$ on the right side of the boulder (moving “downhill”).
We have not attempted to calculate the spatial variation of the drag anisotropy for any actual terrain. Instead, we construct a minimal model that uses simple analytic anisotropy functions $\xi(x)$ that possess the same spatial periodicities as the AA terrain. With $\xi(x)$ in hand, the generalization of the expression for $V_{\rm cm}$ derived in Ref. (8) for small amplitude undulations can be shown to be (see the Supplemental Material)
@@
Fig. 5A plots our RFT results for the average distance traveled by the undulator as a function of $\Gamma/\Delta$ for three values of $\beta$. The good qualitative agreement between the $\beta=(\sqrt{5}-1)/2$ curve (red) and the corresponding experimental data plotted in Fig. 5B (repeated from Fig. 2B) gives us further confidence that RFT captures the essential dynamics of the real robot moving through an AA terrain. The RFT results shown in Fig. 5A for rational values of $\beta$ are surprising (initially) because one reflects ballistic motion ($\beta=1/2$ in green) and the other exhibits localization ($\beta=2/3$ in blue). However, comparison with Fig 5B shows that the robot behaves in exactly the same way. Moreover, the qualitative difference in transport behavior associated with these two rational values of $\beta$ is precisely what is seen in detailed numerical studies of the AA model where the eigenstates of the Schrödinger equation change from delocalized to localized even when $\beta=p/q$ as long as $p\neq 1$ and $q$ are integers(31).
@@
Fig. 6D plots $P_{\rm broad}/P_{\rm 0}$ for the robot calculated similarly using simulation data along with the probability that a trial ends in termination (redrawn from Fig. 2B). Once again, the torque noise rises and falls as a function of the terrain aperiodicity, albeit more steeply in Fig. 6D than in Fig. 6C. The reason for this, we posit, is that the three-dimensional robot can respond to aperiodicity in many ways while the RFT undulator, which is constrained by construction to one-dimension, can do so only by reversals and oscillatory trapping. For example, when $h/d$ is small (below the peak in Fig. 6D), the robot can negotiate the terrain using small out-of-plane motions which produce less broad-band torque noise than a reversal event. As the aperiodicity becomes higher, eventually the robot fails to respond with small out-of-plane motion, and begins to pick up noise in the torque signal. When $h/d$ is large (above the peak), the robot can flip over, jam against a sidewall, or distort into a shape very far from serpenoid. All of the latter events trigger termination, which leads us to conclude that termination simply preempts localization by trapping and likely occurs in response to the same torque patterns. That is why, despite the somewhat ad hoc rules we developed to initiate termination, the systematics of the transition to localization exhibited by our experiments and simulations qualitatively agree with both our RFT calculations and with the behavior expected from the quantum AA model.
@@
The similarity of our results to the cold atom experiments deserves further comment. We are aware that a microscopic quantum particle like an atom and a macroscopic self-deforming object like our robot have nothing in common except that quantum mechanics requires the atom to possess a de Broglie wave, and the act of moving requires our robot to possess a deformation wave. That being said, it is interesting to note (32) that a transition from a delocalized state to a localized state is possible for both classical and quantum particles when they interact in one dimension with the AA potential. Moreover, not unlike our RFT undulator, localization of these particles takes the form of trapping by the wells of the AA potential. The same is not true for the one-dimensional Anderson model (30), where the corresponding potential is random and delocalization can occur for a classical particle but never for a quantum particle.
In closing, we suggest it is worth considering our present results in light of earlier experiments where we observed diffraction-like behavior when snakes and limbless robots passed through a one-dimensional array of posts. We also note work by Zhang and co-workers(40) who found lensing, refraction, and wave-guiding behavior in numerical studies of a filamentary undulator sliding along a heterogeneous surface tailored to mimic optical elements. Taken together, it is tempting to speculate that there is some previously unsuspected connection between the waves associated with self-deforming mechanical systems and the more familiar waves studied in classical and quantum mechanics.
[…]
@@
To complement our experimental investigation of undulatory transport in structured environments, we developed a multibody dynamics simulation using the CoppeliaSim platform (26). The simulation replicates the kinematics of the undulating robot by prescribing a traveling wave of joint angles consistent with a serpenoid gait. While the number of body segments in the simulation was chosen to support computational efficiency and broader parametric exploration, the wavelength of body undulation was kept consistent with that used in physical experiments to preserve comparable locomotor-environment interaction dynamics.
CoppeliaSim was used in conjunction with its Bullet physics (41) backend. Contact interactions were resolved using Bullet’s default sequential impulse constraint solver, an iterative method that applies impulses to enforce non-penetration and friction constraints. This constraint-based formulation enables stable simulation of rigid body interactions in friction-dominated regimes, even in highly cluttered environments.
@@
The robot in simulation was constructed as a planar multibody system using the exact geometry of the physical robot, imported directly from computer-aided design models built in SolidWorks. This ensured that segment dimensions, mass distribution, and contact surface geometry were faithfully preserved, enabling direct comparison between simulated and experimental results. Adjacent segments were connected via revolute joints, allowing relative yaw motion. Joint actuation followed a serpenoid waveform of the form:
@@
Simulations were terminated early if the robot flipped out of the plane, defined as any point on the robot body exceeding a fixed vertical threshold (15 cm, $\leavevmode\nobreak\ \sim 0.9$ width of the body wave) in the $z$-direction. No restitution or damping was applied, allowing motion to cease naturally under frictional contact.
@@
$\frac{dF}{dx}=(C_{N}-C_{L})\frac{dy}{dx}\frac{dy}{dt}-C_{L}\frac{dX}{dt}.$ (7)
The variable $x\in[0,L]$ describes the position along the body in a coordinate system that moves with a point at the tail, while the variable $X(x,t)$ describes the position of a point $x$ along the body in a fixed reference frame as it moves. The variables $C_{N}$ and $C_{L}$ refer to the normal and longitudinal resistance coefficients (in this case, viscous drag coefficients), respectively, which were assumed by Gray and Hancock to be constants intrinsic to the material constituting the body and the environment. In this manuscript, we will eventually allow these constants to vary in space (along $x^{\prime}$), to describe locomotion in heterogeneous environments. We note that this relation is valid in the limit of small displacements $y$.
@@
Substituting (8) into (7) yields
$\frac{dF}{dx}=4\pi^{2}y_{0}^{2}(C_{N}-C_{L})\cos^{2}{2\pi(x+t)}-C_{L}\frac{dX}{dt}$ (9)
@@
We assume that each segment of the body moves along the X-axis at the same speed (the body does not deform along the swimming direction, but only lateral to the swimming direction), such that the integrand on the lhs of (10) is a constant. We now introduce the spatial dependence of the resistance coefficients by writing them as follows
$\frac{C_{N}-C_{L}}{C_{L}}=\frac{C_{N}(X)}{C_{L}(X)}-1=\xi(X)-1$ (11)